As mentioned in the previous post in the week of December 8 many of the staff were working off the facility. This week most everyone is back on site.
The team who went to Texas A&M had a very busy week. They soldered glycol lines, installed covers and tested the system under pressure. They found some leaking at 40 psi and at 80 psi the leaks were severe. The team left it back in the hands of the A&M engineers for them to fix these severe problems. The 2nd VIRUS enclosure is supposed to go on the telescope in January and at that time we would like to do an end to end test of the chilled glycol system.
The team who went to Germany for the HET board of directors meeting brought back good news. The board was pleased to see the progress that the operations team is making and authorized the use of some operations funds for the wide field upgrade to make sure that the teams in Austin and West Texas have enough personnel to stay on schedule for a summer time first light. They were also impressed with the latest installation video, the nitrogen tank installation. Please have a look at: https://www.youtube.com/watch?feature=player_embedded&v=2a1xZF1b0wQ.
The groups that were not traveling have been busy as well. The work on the Anemometer project has drawn to a close. We are collecting data. The next step is to start visualizing the data and actually trying some experiments with louver positions with respect to wind and dome position. This will be a longer term effort.
The night staff have been optimizing the stacks and are now able to get stacks of less than 0.5″. The software team worked this last week to set those optimizations as the default in the stacking code.
An engineer from Austin came out this past week to begin the installation of a clean room for the much anticipated WFU corrector. This huge 4 glass Spherical Aberration corrector will be brought in from University of Arizona on a special truck that will minimize vibration. Even so, when it does arrive it will have to be tested to see if it is still in alignment. This will be done in a clean room tent we are constructing in our loading bay. This will take up the majority of the loading bay for a month and will stop us from receiving any large shipments for the dome floor. This week the frame for the tent went in.
Even over the holiday break we will continue to get some work done. The contractors working on the remote thermal area will be receiving and installing the glycol chiller on site on December 30th. The system probably will not be fully ready to go until mid-January when the 2nd VIRUS enclosure is installed.
Have a look at this nice video the team made of the installation of the enclosure for the HPF:
Since last week included a holiday we will include activities from the last two weeks.
This week we had a software engineer from Austin visiting to commission the Metrology loops in software. The includes the distance measuring interferometer, the tip-tilt sensor and guiding with the CAT. The team was able to accomplish all of these at a single azimuth but problems with the alignment of the CAT stopped further exploration of the performance of the system. We will test the alignment of the CAT on Sunday so that further commissioning can continue.
In the previous blog post we mention the discovery of some loose bolts. The source of the “loosening” is still a mystery but we installed an accelerometer to look for significant jarring events on the tracker. This week we took it down and looked at the data. The most serious events were those of the installation and removal. Even forced Emergency Stops did not produce large accelerations.
Commissioning of the Strip and Wash Room continues. The room heater and make-up air were tested and found to work but needs small tweaking with regards to the indicator lights on the controls. Other members of the team are working on the Safety Protocols document. They believe that commissioning will be completed by January 5th and that we should resume mirror swaps in the primary at that time.
This week one large project has been completed. Phase I of the Dome Sealing and Conditioning was completed with the last of the dome “hair” seals being refurbished. With the completion of Phase I we have sealed the vents on the ring wall, refurbished the seals at the base of the dome, sealed the leaky dome nodes and replaced all of the louver seals. We note that this last rainy season we did not have any leaks in those locations. Phase II involves sealing the dome shutter and will require some engineering.
The technical teams at the HET continue to make progress on various projects such as continuing installing the VCS Safety System components; installing a differential thermostat and a parallel temperature acquisition system in the K-Hut to monitor inside and outside temperatures over the weekend; and test fitting the PFIP Electric Box Frame Weldment Assembly. One pending item that was accomplished was to replace the structure drive wheels. These wheels were showing considerable wear and we replaced them before a failure could occur.
The Anemometer project has made some progress this week. All of the units were calibrated and the system was run over night on Wednesday and Thursday. Some problems were found in shared memory flags in the database but these were cleared up and we should be able to start collecting data and looking at how to best visualize and utilize this new data set to minimize dome seeing induced by turbulence.
In the coming week we will be down a large number of staff. Three of the operations team who are traveling to Texas A&M to speed along the final assembly and testing of the 2nd VIRUS enclosure. Three other people from the operations team (including the blogger) will be traveling to Germany for the HET Board of Director’s Meeting. Much of the last two weeks was spent in preparing reports and presentations for this meeting.